# StimuVAR: Spatiotemporal Stimuli-aware Video Affective Reasoning with Multimodal Large Language Models

### v1

@@ 3.2 Affective Training
To fully integrate the proposed stimuli-aware mechanism into the MLLM backbone and to enhance the MLLM’s affective reasoning ability, we introduce an affective training protocol to fine-tune the MLLM. The affective training consists of two phases. In the initial phase, we train the projector $\mathcal{F}_{proj}$ with training video and emotional response label pairs $\{V,Y_{E}\}$, while keeping the pre-trained visual encoder $\mathcal{F}_{v}$ and LLM $\mathcal{F}_{llm}$ frozen, as follows:
@@
where $\mathcal{L}$ is the cross-entropy loss. This phase aligns the affective visual information with the LLM space and learns the correlation between videos and the emotions they trigger.
In the second phase, we enhance the MLLM’s ability to offer plausible reasoning for its emotional predictions through visual instruction tuning [17, 39, 40]. To create VAR-specific visual instruction data, we use GPT [34, 35] to generate causal connections between videos and the emotions they trigger. Given GPT’s limitation in perceiving large-scale video data, we first employ a vision-language model [53] to produce captions for each sampled frame, ordering them in a format like “Frame 1 description:…; Frame 2 description:…”. GPT then takes temporal correlations into account to summarize these frame-level captions into a video-level caption. Compared to directly captioning videos in a single step using a video-oriented MLLM [19, 20], this progressive summarization method captures video details and ensures frame-level temporal consistency, thereby mitigating hallucination.
Next, inspired by [54], we prompt GPT to generate a reasoning process of deriving the label from an input. Specifically, given the pairs of the video caption <Video Caption> (i.e., the caption of $V$) and the emotional response label <Emotion> (i.e., $Y_{E}$), we query GPT to explain viewers’ emotional responses when they watch the videos. An example of our prompt is shown below: {‘‘role’’: ‘‘system’’, ‘‘content’’: Given the below (QUESTION, ANSWER) pair examples of emotion estimation, left fill-in the REASONING process which derives ANSWERS from QUESTIONS in three sentences.}, {‘‘role’’: ‘‘user’’, ‘‘content’’: QUESTION: These are frame descriptions from a video. After reading the descriptions, how people might emotionally feel about the content and why. Only provide the one most likely emotion. <Video Caption> ANSWER: The viewer feels <Emotion>. REASONING: Lets think of step-by-step The output reasoning process is denoted as $Y_{R}$. We use our VAR instruction data $\{V,Y_{E},Y_{R}\}$ to train the emotion-triggered tube module $\theta_{tube}$ and fine-tune the projector and LLM (by LoRA [55]) as follows:
@@
where $\mathcal{L}$ is the cross-entropy loss. This phase fosters the correspondence between visual tokens and output text, enhancing the MLLM’s affective reasoning ability. Our affective training protocol directs the MLLM’s reasoning strengths and commonsense knowledge towards an emotional focus, enabling the MLLM to offer plausible explanations for its affective understanding.
@@ 4.1 Experimental Setup
Datasets. We evaluate our method on two viewer-centered VEA datasets. Video Cognitive Empathy (VCE) [13] contains 61,046 videos with 27 fine-grained emotional responses [56] manually annotated. The dataset is divided into 50,000 videos for training and 11,046 for testing, with each video averaging 14.1 seconds in length. This dataset purposely removes audio cues to focus solely on visual information. VideoEmotion-8 (VE-8) [14] comprises 1,101 user-generated videos sourced from YouTube and Flickr. Each video is categorized into one of eight emotions from Plutchik’s Wheel of Emotions [57] and has an average duration of around 107 seconds.
Evaluation metrics. We follow the official metrics used by the VCE and VE-8 datasets. VCE utilizes the Top-3 accuracy, while VE-8 evaluates the Top-1 accuracy. Additionally, we provide a comprehensive protocol with extensive metrics to evaluate the reasoning quality for the VAR problem. All the metrics are reference-free since existing viewer-centered VEA datasets [13, 14] do not contain ground-truth reasoning. These metrics are described as follows:
@@
    Emotional-alignment (Emo-align) [11] uses GPT-3.5 [34] as a text-to-emotion classifier $C_{E|R}$ that measures the accuracy (%) of the emotion predicted from reasoning text. This metric evaluates if the output reasoning $R$ evokes the same emotion $E$ as the ground truth $Y_{E}$.
@@
    Doubly-right {RR, RW, WR, WW[36] jointly considers the accuracy of prediction and reasoning, where RR denotes the percentage (%) of Right emotional prediction with Right reasoning, RW denotes Right prediction with Wrong reasoning, WR denotes Wrong prediction with Right reasoning, and WW denotes Wrong prediction with Wrong reasoning. Here the correctness of reasoning is defined by Emo-align. We desire a high accuracy of RR (the best is 100%) and low percentages of RW, WR and WW (the best is 0%). The sum of {RR, RW, WR, WW} is 100%.
@@
    CLIPScore (CLIP-S) [37] uses CLIP [58] to measure the compatibility between input videos and output reasoning in feature space. A higher score indicates better video-reasoning compatibility.
@@
    LLM-as-a-judge [38] uses GPT-3.5 as a judge to rate the quality of reasoning. We set the score as an integer within a range of $[1,4]$. In addition to average scores, we also report the 1-vs.-1 (StimuVAR vs. a baseline) comparison results showing the number of test data where StimuVAR wins, loses or ties.
Baseline approaches. To our knowledge, StimuVAR may be the first method for the viewer-centered VAR problem. Hence, we broadly compare it with traditional zero-shot approaches [13, 58], traditional emotion models [59, 60, 61, 16], and state-of-the-art video-oriented MLLMs [19, 20, 21, 22, 23, 24, 25]. The traditional emotion models are specifically trained for emotion tasks, but they cannot offer the rationales behind their predictions. We also consider EmoVIT [17], a state-of-the-art emotion MLLM for image data, and we extend it to videos for comparison.
Implementation details. The MLLM backbone uses CLIP ViT-L/14 [58, 62] as the visual encoder, Llama2-7b [42] as the LLM, and a linear layer as the projector. For VAR instruction data creation, we employ CogVLM-17B [53] to generate captions for each selected frame and GPT-4 [35] to aggregate these frame-level captions into a video caption. GPT-3.5 [34] is then used to generate causal connections between the videos and the emotions they trigger. In event-driven frame sampling, we set $N=6$ as the default number of sampled frames. The emotion-triggered tube selection module refines the <Patch> tokens after the projector, retaining the top $K=4$ tubes with shape $2\times 4\times 4$ as spatial tokens and <CLS> as temporal tokens to represent a video. The model is trained using the AdamW optimizer [63] with a learning rate of $5\times 10^{-5}$ and a batch size of $1$, through $2$ and $3$ epochs in the phase I and II of affective training, respectively.
@@
CLIP [58] ICML’21 28.4 - - - - - -
Majority [13] NeurIPS’22 35.7 - - - - - -
R(2+1)D [59] CVPR’18 65.6 - - - - - -
STAM [60] arXiv’21 66.4 - - - - - -
VideoMAE [61] NeurIPS’22 68.9 - - - - - -
MM-VEMA [16] PRCV’23 73.3 - - - - - -
@@
Video-LLaMA [19] EMNLP’23 26.4 25.5 16.2 10.2 9.3 64.3 63.9
Video-LLaVA [20] arXiv’23 25.0 31.2 17.5 7.5 13.7 61.3 70.6
Valley [21] arXiv’23 31.3 29.4 19.2 12.1 10.2 58.5 69.4
Video-ChatGPT [22] ACL’24 21.0 29.5 11.4 9.5 18.1 61.0 68.9
VideoChat2 [23] CVPR’24 31.1 36.4 24.0 7.1 12.4 56.5 68.6
Chat-UniVi [24] CVPR’24 38.6 29.5 21.0 17.6 8.5 52.9 70.2
mPLUG-Owl [25] CVPR’24 23.6 22.1 13.8 9.7 8.3 68.2 69.3
EmoVIT [17] CVPR’24 10.5 5.2 4.8 5.7 0.4 89.1 48.9
@@
Video-LLaMA [19] EMNLP’23 12.4 21.6 10.4 1.9 11.3 76.4 62.3
Video-LLaVA [20] arXiv’23 21.5 20.5 17.7 3.8 2.8 75.7 61.3
Valley [21] arXiv’23 16.1 20.6 12.2 4.0 8.5 75.3 67.0
Video-ChatGPT [22] ACL’24 17.2 16.6 14.1 3.1 2.5 80.2 66.0
VideoChat2 [23] CVPR’24 16.5 15.6 14.3 2.2 1.4 82.1 65.1
Chat-UniVi [24] CVPR’24 14.4 19.2 9.9 4.5 9.3 76.3 65.9
mPLUG-Owl [25] CVPR’24 26.2 24.7 19.1 7.0 5.6 68.3 64.4
EmoVIT [17] CVPR’24 25.1 0.3 0.3 24.8 0.0 74.9 45.9
@@
Datasets VCE [13]      VE-8 [14]
Judgement Win Lose Tie Score Win Lose Tie Score
Video-LLaMA [19] 41 11 48 3.03 48 13 39 2.83
Video-LLaVA [20] 45 13 42 3.02 49 23 28 2.45
Valley [21] 45 12 43 2.89 35 7 58 2.88
Video-ChatGPT [22] 54 8 38 2.69 30 18 52 3.12
VideoChat2 [23] 56 12 32 2.65 48 13 39 3.02
Chat-UniVi [24] 50 19 31 2.83 37 20 43 3.08
mPLUG-Owl [25] 37 17 46 3.18 29 20 51 3.14
EmoVIT [17] 79 2 19 2.18 92 0 8 1.99
StimuVAR (Ours) - - - 3.40 - - - 3.29
@@
4.2 Main Results
Table 1 reports the results on the VCE dataset. Traditional emotion models generally outperform MLLMs in prediction accuracy since they are specifically trained for emotion tasks. However, these traditional models are not able to provide explanations for their predictions. In contrast, MLLMs have a significant reasoning capability. Compared to existing video-oriented MLLMs, the proposed StimuVAR demonstrates superiority in the VAR task, both in predicting viewers’ emotional responses and offering rationales behind its predictions. We can observe that StimuVAR achieves 69.6% in Emo-align, which is double the performance of most MLLM baselines. It also has 68.8% in RR, which shows consistency between emotional predictions and reasoning. StimuVAR’s highest CLIP-S score indicates a better correspondence between input videos and output reasoning, suggesting that its reasoning closely adheres to the visual cues and covers details. Notably, while EmoVIT [17] is an MLLM specifically trained with emotion knowledge, it does not perform well on viewer-centered VEA datasets. The reason may be that EmoVIT prioritizes image attributes, whereas video data require temporal attention. This validates the importance of a spatiotemporal stimuli-aware mechanism. Table 3 presents the LLM-as-a-judge results. Due to GPT’s query limitation, we randomly select 100 test videos for this metric. StimuVAR achieves the highest GPT rating scores and better winning percentages against all competitors in 1-vs.-1 comparisons, demonstrating its coherent and insightful explanations.
Table 2 and Table 3 present the results on the VE-8 dataset. To test the generalizability of StimuVAR, we evaluate the model trained with VCE data without fine-tuning on VE-8. Although VE-8 has a different emotion categorization, lower video quality, and nearly eight times longer video duration, StimuVAR again outperforms all the MLLM baselines. Furthermore, this performance is achieved by sampling only six frames, making the computational efficiency more significant for these long videos. These results demonstrate that the proposed emotional stimuli awareness has high generalizability and can efficiently reduce computational costs. Overall, StimuVAR consistently achieves the best scores across all the metrics on both datasets.

### v2

@@ 3.2 Affective Training
To fully integrate the proposed stimuli-aware mechanism into the MLLM backbone and to enhance the MLLM’s affective reasoning ability, we introduce an affective training protocol to fine-tune the MLLM.
3.2.1 VAR Visual Instruction Data Construction
To prepare effective affective training, we construct VAR-specific visual instruction data based on the training set of the raw Video Cognitive Empathy (VCE) dataset [mazeika2022would], a large-scale viewer-centered video dataset with fine-grained annotations. Detailed dataset information is provided in Sec. 4.1. Given the strong capabilities of GPT [brown2020language, achiam2023gpt] for constructing logical reasoning across concepts, we leverage it to generate instructions with causal relationships between videos and the emotions they evoke.
Due to GPT’s limitation in directly processing large-scale video data, we first employ a vision-language model[wang2023cogvlm] to generate captions for each sampled frame, structuring them sequentially in a format of “Frame 1 description: …; Frame 2 description: …”. GPT then processes these frame-level captions, incorporating temporal correlations to produce a coherent video-level caption. Compared to directly captioning videos in a single step using a video-oriented MLLM [Zhang2023VideoLLaMAAI, Lin2023VideoLLaVALU], this progressive summarization method captures video details and ensures frame-level temporal consistency, thereby mitigating hallucination.
Next, inspired by [zhang2024mm], we prompt GPT to generate a reasoning process of deriving the label from an input. Specifically, given the pairs of the video caption <Video Caption> (i.e., the caption of $V$) and the emotional response label <Emotion> (i.e., $Y_{E}$), we query GPT to explain viewers’ emotional responses when they watch the videos. This approach enables us to generate detailed video captions along with a structured rationale that explicitly connects viewers’ emotional responses to the underlying stimuli. Examples of video captions and instructions are provided in Supplementary Material. An example of our prompt is shown below, and its output reasoning is denoted as $Y_{R}$:
{‘‘role’’: ‘‘system’’,
‘‘content’’: Given the below (QUESTION, ANSWER) pair examples of emotion estimation, left fill-in the REASONING process which derives ANSWERS from QUESTIONS in three sentences.},
{‘‘role’’: ‘‘user’’,
‘‘content’’: QUESTION: These are frame descriptions from a video. After reading the descriptions, how people might emotionally feel about the content and why. Only provide the one most likely emotion. <Video Caption>
ANSWER: The viewer feels <Emotion>.
REASONING: Lets think of step-by-step
3.2.2 Two-phase Training
Our affective training consists of two phases (see Fig. 2). In the initial phase, we use the original training set of the VCE dataset with videos and emotional response label pairs $\{V,Y_{E}\}$, where only the projector $\theta_{\mathcal{F}_{proj}}$ of the MLLM is trainable, while keeping the pre-trained visual encoder $\mathcal{F}_{v}$ and LLM $\mathcal{F}_{llm}$ frozen, as follows:
@@ 3.2.2 Two-phase Training
where $\mathcal{L}$ is the cross-entropy loss. During this phase, the projector aligns the visual features extracted by the pre-trained visual encoder with the pre-trained LLM word embeddings and learns the correlation between videos and the emotions they trigger, supervised by the ground-truth emotion annotations $Y_{E}$.
In the second phase, we enhance the MLLM’s ability to offer plausible reasoning for its emotional predictions through visual instruction tuning [xie2024emovit, liu2024visual, Dai2023InstructBLIPTG]. We use our VAR instruction data $\{V,Y_{E},Y_{R}\}$ (based on the original training set of the VCE dataset $\{V,Y_{E}\}$) constructed in Sec. 3.2.1 to train the proposed emotion-triggered tube selection module $\theta_{tube}$ and fine-tune the projector $\theta_{\mathcal{F}_{proj}}$ and the LLM (using LoRA [hu2022lora]) $\theta^{lora}_{\mathcal{F}_{llm}}$. The training objective is formulated as follows:
@@
where $\mathcal{L}$ is the cross-entropy loss. This fine-tuning enhances the model’s affective reasoning capabilities for its emotional predictions. Furthermore, the emotion-triggered tube selection module strengthens the link between visual tokens of stimuli and the elements referenced in the reasoning, while also filtering out irrelevant visual tokens, reducing noise, improving decision-making, and enhancing the model’s efficiency.
Our two-phase affective training protocol directs the MLLM’s reasoning strengths and commonsense knowledge towards an emotional focus, enabling the MLLM to offer plausible explanations for its affective understanding.
@@ 4.1 Experimental Setup
Datasets. We evaluate our method on three viewer-centered VEA datasets. Video Cognitive Empathy (VCE) [mazeika2022would] contains 61,046 videos with 27 fine-grained emotional responses [cowen2017self] manually annotated. The dataset is divided into 50,000 videos for training and 11,046 for testing, with each video averaging 14.1 seconds in length. This dataset purposely removes audio cues to focus solely on visual information. VideoEmotion-8 (VE-8) [jiang2014predicting] comprises 1,101 user-generated videos sourced from YouTube and Flickr. Each video is categorized into one of eight emotions from Plutchik’s Wheel of Emotions [plutchik1980general] and has an average duration of around 107 seconds. YouTube/Flickr-EkmanSix (YF-6) [xu2016heterogeneous] shares the same video source as VE-8 but expands the size to 1,637, categorized into Ekman’s six basic emotions [ekman1971constants]. Since our VAR instruction data is constructed using the training set of VCE, we employ the VCE test set for in-domain evaluation while using VE-8 and YF-6 as out-of-domain datasets to assess our method’s effectiveness and generalization capabilities.
Evaluation metrics. We follow the official metrics used by the VCE, VE-8 and YF-6 datasets. VCE utilizes the Top-3 accuracy, while VE-8 and YF-6 evaluate the Top-1 accuracy. Additionally, we provide a comprehensive protocol with extensive metrics to evaluate the reasoning quality for the VAR problem. All the metrics are reference-free since existing viewer-centered VEA datasets [mazeika2022would, jiang2014predicting] do not contain ground-truth reasoning. These metrics are described as follows:
@@
    Emotional-alignment (Emo-align) [achlioptas2021artemis] uses GPT-3.5 [brown2020language] as a text-to-emotion classifier $C_{E|R}$ that measures the accuracy (%) of the emotion predicted from reasoning text. This metric evaluates if the output reasoning $R$ evokes the same emotion $E$ as the ground truth $Y_{E}$.
@@
    Doubly-right {RR, RW, WR, WW[mao2023doubly] jointly considers the accuracy of prediction and reasoning, where RR denotes the percentage (%) of Right emotional prediction with Right reasoning, RW denotes Right prediction with Wrong reasoning, WR denotes Wrong prediction with Right reasoning, and WW denotes Wrong prediction with Wrong reasoning. Here the correctness of reasoning is defined by Emo-align. We desire a high accuracy of RR (the best is 100%) and low percentages of RW, WR and WW (the best is 0%). The sum of {RR, RW, WR, WW} is 100%.
@@
    CLIPScore (CLIP-S) [hessel2021clipscore] uses CLIP [radford2021learning] to measure the compatibility between input videos and output reasoning in feature space. A higher score indicates better video-reasoning compatibility.
@@
    LLM-as-a-judge [zheng2024judging] uses GPT-3.5 [brown2020language] and Claude 3.5 [anthropic2024claude] as a judge respectively to rate the quality of reasoning. We set the score as an integer within a range of $[1,4]$. In addition to average scores, we also report the 1-vs.-1 (StimuVAR vs. a baseline) comparison results showing the number of test data where StimuVAR wins, loses or ties.
Baseline approaches. To our knowledge, StimuVAR may be the first method for the viewer-centered VAR problem. Hence, we broadly compare it with traditional zero-shot approaches [mazeika2022would, radford2021learning], traditional emotion models [tran2018closer, sharir2021image, tong2022videomae, pu2023going], state-of-the-art video-oriented MLLMs [Zhang2023VideoLLaMAAI, Lin2023VideoLLaVALU, luo2023valley, Maaz2023VideoChatGPTTD, li2024mvbench, Jin_2024_CVPR, Ye_2024_CVPR], and a long-video MLLM, LongVLM [weng2024longvlm]. The traditional emotion models are specifically trained for emotion tasks, but they cannot offer the rationales behind their predictions. We also consider EmoVIT [xie2024emovit], a state-of-the-art emotion MLLM for image data, and we extend it to videos for comparison.
Implementation details. The MLLM backbone uses CLIP ViT-L/14 [radford2021learning, dosovitskiy2021image] as the visual encoder, Llama2-7b [touvron2023llama] as the LLM, and a linear layer as the projector. For VAR instruction data creation, we employ CogVLM-17B [wang2023cogvlm] to generate captions for each selected frame and GPT-4 [achiam2023gpt] to aggregate these frame-level captions into a video caption. GPT-3.5 [brown2020language] is then used to generate causal connections between the videos and the emotions they trigger. In event-driven frame sampling, we set $N=6$ as the default number of sampled frames. The emotion-triggered tube selection module refines the <Patch> tokens after the projector, retaining the top $K=4$ tubes with shape $2\times 4\times 4$ as spatial tokens and <CLS> as temporal tokens to represent a video. The model is trained using the AdamW optimizer [loshchilov2019decoupled] with a learning rate of $5\times 10^{-5}$ and a batch size of $1$, through $2$ and $3$ epochs in the phase I and II of affective training, respectively.
@@
CLIP [radford2021learning] ICML’21 28.4 - - - - - -
Majority [mazeika2022would] NeurIPS’22 35.7 - - - - - -
R(2+1)D [tran2018closer] CVPR’18 65.6 - - - - - -
STAM [sharir2021image] arXiv’21 66.4 - - - - - -
VideoMAE [tong2022videomae] NeurIPS’22 68.9 - - - - - -
MM-VEMA [pu2023going] PRCV’23 73.3 - - - - - -
@@
Video-LLaMA [Zhang2023VideoLLaMAAI] EMNLP’23 26.4 25.5 16.2 10.2 9.3 64.3 63.9
Video-LLaVA [Lin2023VideoLLaVALU] arXiv’23 25.0 31.2 17.5 7.5 13.7 61.3 70.6
Valley [luo2023valley] arXiv’23 31.3 29.4 19.2 12.1 10.2 58.5 69.4
Video-ChatGPT [Maaz2023VideoChatGPTTD] ACL’24 21.0 29.5 11.4 9.5 18.1 61.0 68.9
VideoChat2 [li2024mvbench] CVPR’24 31.1 36.4 24.0 7.1 12.4 56.5 68.6
Chat-UniVi [Jin_2024_CVPR] CVPR’24 38.6 29.5 21.0 17.6 8.5 52.9 70.2
mPLUG-Owl [Ye_2024_CVPR] CVPR’24 23.6 22.1 13.8 9.7 8.3 68.2 69.3
EmoVIT [xie2024emovit] CVPR’24 10.5 5.2 4.8 5.7 0.4 89.1 48.9
LongVLM [weng2024longvlm] ECCV’24 20.9 19.7 6.9 13.9 12.8 66.4 64.3
@@
Datasets VCE [mazeika2022would]      VE-8 [jiang2014predicting]
Judgement Win Lose Tie Score Win Lose Tie Score
Video-LLaMA [Zhang2023VideoLLaMAAI] 41 11 48 3.03 48 13 39 2.83
Video-LLaVA [Lin2023VideoLLaVALU] 45 13 42 3.02 49 23 28 2.45
Valley [luo2023valley] 45 12 43 2.89 35 7 58 2.88
Video-ChatGPT [Maaz2023VideoChatGPTTD] 54 8 38 2.69 30 18 52 3.12
VideoChat2 [li2024mvbench] 56 12 32 2.65 48 13 39 3.02
Chat-UniVi [Jin_2024_CVPR] 50 19 31 2.83 37 20 43 3.08
mPLUG-Owl [Ye_2024_CVPR] 37 17 46 3.18 29 20 51 3.14
EmoVIT [xie2024emovit] 79 2 19 2.18 92 0 8 1.99
StimuVAR (Ours) - - - 3.40 - - - 3.29
4.2 Main Results
Table 1 reports the results on the VCE dataset. Traditional emotion models generally outperform MLLMs in prediction accuracy since they are specifically trained for emotion tasks. However, these traditional models are not able to provide explanations for their predictions. In contrast, MLLMs have a significant reasoning capability. Compared to existing video-oriented MLLMs, the proposed StimuVAR demonstrates superiority in the VAR task, both in predicting viewers’ emotional responses and offering rationales behind its predictions. We can observe that StimuVAR achieves 69.6% in Emo-align, which is double the performance of most MLLM baselines. It also has 68.8% in RR, which shows consistency between emotional predictions and reasoning. StimuVAR’s highest CLIP-S score indicates a better correspondence between input videos and output reasoning, suggesting that its reasoning closely adheres to the visual cues and covers details. Notably, while EmoVIT [xie2024emovit] is an MLLM specifically trained with emotion knowledge, it does not perform well on viewer-centered VEA datasets. The reason may be that EmoVIT prioritizes image attributes, whereas video data require temporal attention. This validates the importance of a spatiotemporal stimuli-aware mechanism. Table 2 presents the LLM-as-a-judge results using GPT. Due to GPT’s query limitation, we randomly select 100 test videos for this metric. StimuVAR achieves the highest GPT rating scores and better winning percentages against all competitors in 1-vs.-1 comparisons, demonstrating its coherent and insightful explanations. The results using Claude 3.5 as a judge are reported in the Supplementary Material.
Table 2 and Table 3 present the results on the VE-8 dataset. To test the generalizability of StimuVAR, we evaluate the model trained with VCE data without fine-tuning on VE-8. Although VE-8 has a different emotion categorization, lower video quality, and nearly eight times longer video duration, StimuVAR again outperforms all the MLLM baselines. Furthermore, this performance is achieved by sampling only six frames, making the computational efficiency more significant for these long videos. Table 4 shows a performance ranking in the YF-6 dataset similar to that of VE-8, where StimuVAR achieves the highest Top-1 accuracy of 46.2%. These results demonstrate that the proposed emotional stimuli awareness has high generalizability and can efficiently reduce computational costs. Overall, StimuVAR consistently achieves the best scores across all the metrics on both datasets.
@@ 4.2 Main Results
Video-LLaMA [Zhang2023VideoLLaMAAI] EMNLP’23 12.4 21.6 10.4 1.9 11.3 76.4 62.3
Video-LLaVA [Lin2023VideoLLaVALU] arXiv’23 21.5 20.5 17.7 3.8 2.8 75.7 61.3
Valley [luo2023valley] arXiv’23 16.1 20.6 12.2 4.0 8.5 75.3 67.0
Video-ChatGPT [Maaz2023VideoChatGPTTD] ACL’24 17.2 16.6 14.1 3.1 2.5 80.2 66.0
VideoChat2 [li2024mvbench] CVPR’24 16.5 15.6 14.3 2.2 1.4 82.1 65.1
Chat-UniVi [Jin_2024_CVPR] CVPR’24 14.4 19.2 9.9 4.5 9.3 76.3 65.9
mPLUG-Owl [Ye_2024_CVPR] CVPR’24 26.2 24.7 19.1 7.0 5.6 68.3 64.4
EmoVIT [xie2024emovit] CVPR’24 25.1 0.3 0.3 24.8 0.0 74.9 45.9
@@
Table 4: Quantitative comparison on the YF-6 dataset.
Method Acc Emo-ali RR RW WR WW CLIP-S Win Lose Tie Score
Video-LLaMA [Zhang2023VideoLLaMAAI] 17.3 17.7 10.5 6.8 7.2 75.5 62.4 44 12 44 2.80
Video-LLaVA [Lin2023VideoLLaVALU] 20.1 17.5 17.1 3.0 0.4 79.5 66.8 42 17 41 2.85
Valley [luo2023valley] 10.0 18.9 5.9 4.1 13.0 77.0 66.4 44 10 46 2.89
Video-ChatGPT [Maaz2023VideoChatGPTTD] 17.9 19.7 13.2 4.7 6.5 75.6 63.8 43 16 41 2.93
VideoChat2 [li2024mvbench] 24.3 23.7 18.5 5.8 5.2 70.5 64.1 37 16 47 3.03
Chat-UniVi [Jin_2024_CVPR] 28.4 29.2 22.1 6.3 7.1 64.5 66.3 33 14 53 3.12
mPLUG-Owl [Ye_2024_CVPR] 22.8 24.9 14.9 7.9 10.0 67.2 53.7 31 17 52 3.11
EmoVIT [xie2024emovit] 0.7 1.7 0.5 0.2 1.2 98.1 52.5 85 0 15 1.88
StimuVAR (Ours) 46.2 44.2 37.3 8.9 6.9 46.9 72.1 - - - 3.30